if 0.12mol of zinc reacts with excess lead (iv) sulfate, how many grams of zinc sulfate will be produced in the reaction
1
Expert's answer
2019-10-14T14:29:53-0400

2Zn+Pb(SO4)2->2ZnSO4+Pb

n(ZnSO4)=n(Zn)=0.12mol

m(Zn)=M(Zn)*n(Zn)=65,38*0.12=7.85g